# Spiking Neural Network as Adaptive Event Stream Slicer

### v1

@@ Abstract
Event-based cameras are attracting significant interest as they provide rich edge information, high dynamic range, and high temporal resolution. Many state-of-the-art event-based algorithms rely on splitting the events into fixed groups, resulting in the omission of crucial temporal information, particularly when dealing with diverse motion scenarios (e.g., high/low speed). In this work, we propose SpikeSlicer, a novel-designed plug-and-play event processing method capable of splitting events stream adaptively. SpikeSlicer utilizes a lightweight (0.41M) and low-energy spiking neural network (SNN) to trigger event slicing. To guide the SNN to fire spikes at optimal time steps, we propose the Spiking Position-aware Loss (SPA-Loss) to modulate the neuron’s state. Additionally, we develop a Feedback-Update training strategy that refines the slicing decisions using feedback from the downstream artificial neural network (ANN). Extensive experiments demonstrate that our method yields significant performance improvements in event-based object tracking and recognition. Notably, SpikeSlicer provides a brand-new SNN-ANN cooperation paradigm, where the SNN acts as an efficient, low-energy data processor to assist the ANN in improving downstream performance, injecting new perspectives and potential avenues of exploration.
@@ 1 Introduction
  • Extensive experiments demonstrate that SpikeSlicer significantly improves model performance by up to 22.3% in object tracking and 19.2% in recognition with only a 2.7% increase in energy consumption.
@@ 3.2 Adaptive Event Slicing Process
Considering an event stream $\mathcal{E}$, we first convert $\mathcal{E}$ into a list of time-continuous event cells. Event cells are then continuously entered into a lightweight SNN (structure details are provided in Appendix K) through a loop operation. Through forward propagation, the features of the last hidden layers ($h^{L-1}$) are finally mapped to a single spiking neuron to activate spikes:
@@ 4.1 Beginner’s Arena: Event Slicing in Simple Tasks
We compare the SPA-Loss function with common MSE-Loss and CE-Loss. The experiment setting is detailed in Appendix J. As depicted in Fig 4(b), the SPA-Loss successfully supervises the SNN to activate spikes at the desired time steps. In particular, SPA-Loss requires only a few iterations ($<$400) to supervise the SNN to fire spikes correctly. In contrast, MSE-Loss can only succeed at certain time steps, and CE-Loss cannot even accomplish the task. In addition, using both Mem-Loss and LA-Loss yields smoother results compared to using Mem-Loss alone. To summarize, the beginner’s arena preliminarily tests the effectiveness of the SPA-Loss and paves the way for subsequent experiments.
@@ 4.2 Expert’s Arena: Mastering Adaptive Event Slicing with SNN-ANN Collaboration
After a successful challenge in the beginner’s arena, we move on to the expert arena. Here we use a lightweight SNN (0.41M) to adaptively process the event data on complex downstream tasks:
@@ 4.3 Analysis of the Adaptive Slicing Method
Analysis of Efficiency and Speed. We evaluate the performance of the ANN tracker (TransT) with or without the SpikeSlicer in terms of energy consumption and processing speed. Detailed calculations for energy consumption are in Appendix L. In Tab. 3, the SpikeSlicer stands out as a lightweight, plug-and-play option, adding only 0.41M parameters to the system. The introduction of the SpikeSlicer incurs a marginal energy increase of 2.7% and a 14.1% reduction in processing speed. However, these costs are offset by a significant 22.3% improvement in the RSR metric.
Table 3: Comparison of Efficiency and Speed. The comparison includes the number of operations (OPs) and frames per second (FPS), where FPS values are derived from GPU processing speeds.
Models # Parameter (M) OPs (G) Energy (mJ) FPS (Hz) Performance
ANN w/o SpikeSlicer 23.01 56.36 259.26 149.25 51.0
ANN w/ SpikeSlicer 23.42 61.01 266.31 128.21 62.4
@@ 4.4 Ablation Study
Ablation Study on Event Cell Number. Considering that the number of event cells $N$ may affect the SNN’s decision on event slicing, we examine the stability of the sliced event group by varying the size of $N$. Tab. 6 shows that the average spike time of the SNN varies for different $N$, but the time duration (i.e., Avg $\text{Avg Spike}/N$) of the resulting event groups is stable. This verifies that the SNN can effectively make cuts based on event information rather than making decisions based on the number of inputs alone. More details are provided in Appendix O.
@@ 5 Limitation
We summarize the limitations of this work as follows: Firstly, the SpikeSlicer process involves multi-stage SNN-ANN training, which leaves substantial room for improvement in the adaptive slicing strategy. Secondly, for recognition tasks, we convert the event stream into a single-frame representation to obtain accurate supervisory signals. This approach could be refined in the future to enable SpikeSlicer to slice stream events into multi-frame representations, which are the mainstream format. Thirdly, our experiments are conducted on GPUs; however, the most suitable hardware for SNNs would be brain-inspired chips. As a result, conducting SNN experiments on GPUs may lead to slower overall inference speeds. Extending this paradigm to brain-inspired chips is an interesting direction worth exploring in the future.
@@ I.3 Details of Dynamic Hyperparameter Tuning
Appendix J Implementation Details in Toy Experiments
@@
In detail, we set the time step within the range $[1,30]$ and the max number of iterations as 800. We utilize a lightweight convolutional SNN with random initialization. We compare our proposed SPA-Loss function with common mean square error (MSE-Loss) and cross-entropy loss (CE-Loss). The results in Fig. 4 demonstrate that our proposed SPA-Loss can guide the SNN to spike at the desired timestep accurately with fewer convergence iterations than standard MSE-Loss and CE-Loss, paving the way for complex event-based vision tasks in Sec. 4.2. More experiments on the beginner’s arena are shown in Appendix M.
Appendix K Implementation Details in Event-based Task
@@
K.1 Event-based Object Tracking
@@
K.2 Event-based Recognition
@@
Appendix L Theoretical Energy Consumption Calculation
@@
Appendix M More Experiments on Beginner’s Arena
@@
Appendix N Statistics of Dynamic Slicing (SpikeSlicer) vs. Fixed Slicing
@@
Appendix O Statistics of Resulting Slicing Duration
@@
Appendix P More Experiments with Latest Models
@@
Appendix Q More Experiments with Complex Neuromorphic Dataset
@@
Appendix R Impact Statement

### v2

@@ Abstract
Event-based cameras are attracting significant interest as they provide rich edge information, high dynamic range, and high temporal resolution. Many state-of-the-art event-based algorithms rely on splitting the events into fixed groups, resulting in the omission of crucial temporal information, particularly when dealing with diverse motion scenarios (e.g., high/low speed). In this work, we propose SpikeSlicer, a novel-designed plug-and-play event processing method capable of splitting events stream adaptively. SpikeSlicer utilizes a low-energy spiking neural network (SNN) to trigger event slicing. To guide the SNN to fire spikes at optimal time steps, we propose the Spiking Position-aware Loss (SPA-Loss) to modulate the neuron’s state. Additionally, we develop a Feedback-Update training strategy that refines the slicing decisions using feedback from the downstream artificial neural network (ANN). Extensive experiments demonstrate that our method yields significant performance improvements in event-based object tracking and recognition. Notably, SpikeSlicer provides a brand-new SNN-ANN cooperation paradigm, where the SNN acts as an efficient, low-energy data processor to assist the ANN in improving downstream performance, injecting new perspectives and potential avenues of exploration.
@@ 1 Introduction
  • Extensive experiments demonstrate that SpikeSlicer significantly improves model performance by up to 22.3% in object tracking and 19.2% in recognition with only a 0.32% increase in energy consumption.
@@ 3.2 Adaptive Event Slicing Process
Considering an event stream $\mathcal{E}$, we first convert $\mathcal{E}$ into a list of time-continuous event cells. Event cells are then continuously entered into a SNN (structure details are provided in Appendix L) through a loop operation. Through forward propagation, the features of the last hidden layers ($h^{L-1}$) are finally mapped to a single spiking neuron to activate spikes:
@@ 4.1 Beginner’s Arena: Event Slicing in Simple Tasks
We compare the SPA-Loss function with common MSE-Loss and CE-Loss. The experiment setting is detailed in Appendix K. As depicted in Fig 4(b), the SPA-Loss successfully supervises the SNN to activate spikes at the desired time steps. In particular, SPA-Loss requires only a few iterations ($<$400) to supervise the SNN to fire spikes correctly. In contrast, MSE-Loss can only succeed at certain time steps, and CE-Loss cannot even accomplish the task. In addition, using both Mem-Loss and LA-Loss yields smoother results compared to using Mem-Loss alone. To summarize, the beginner’s arena preliminarily tests the effectiveness of the SPA-Loss and paves the way for subsequent experiments.
@@ 4.2 Expert’s Arena: Mastering Adaptive Event Slicing with SNN-ANN Collaboration
After a successful challenge in the beginner’s arena, we move on to the expert arena. Here we use a low-energy SNN to adaptively process the event data on complex downstream tasks:
@@ 4.3 Analysis of the Adaptive Slicing Method
Analysis of Efficiency and Speed. We evaluate the performance of the ANN tracker (TransT) with or without the SpikeSlicer in terms of energy consumption and processing speed. Detailed calculations for energy consumption are in Appendix M. In Tab. 3, the SpikeSlicer stands out as a plug-and-play option. The introduction of the SpikeSlicer incurs a marginal energy increase of 0.85 mJ and a small reduction in processing speed. However, these costs are offset by a significant 22.3% improvement in the RSR metric.
Table 3: Comparison of Efficiency and Speed. The comparison includes the number of operations (OPs) and tracking speed per image without image processing time.
Models OPs (G) Energy (mJ) Speed (s / img) Performance
ANN w/o SpikeSlicer 56.36 259.26 0.045 51.0
ANN w/ SpikeSlicer 57.09 260.11 0.060 62.4
@@ 4.4 Ablation Study
Ablation Study on Event Cell Number. Considering that the number of event cells $N$ may affect the SNN’s decision on event slicing, we examine the stability of the sliced event group by varying the size of $N$. Tab. 6 shows that the average spike time of the SNN varies for different $N$, but the time duration (i.e., Avg $\text{Avg Spike}/N$) of the resulting event groups is stable. This verifies that the SNN can effectively make cuts based on event information rather than making decisions based on the number of inputs alone. More details are provided in Appendix P.
@@ 5 Limitation
We summarize the limitations of this work as follows: Firstly, the SpikeSlicer process involves multi-stage SNN-ANN training, which leaves substantial room for improvement in the adaptive slicing strategy. Secondly, for recognition tasks, we convert the event stream into a single-frame representation to obtain accurate supervisory signals. This approach could be refined in the future to enable SpikeSlicer to slice stream events into multi-frame representations, which are the mainstream format. Thirdly, our experiments are conducted on GPUs; however, the most suitable hardware for SNNs would be brain-inspired chips. In addition, dynamic events need to be generated and processed in real-time during inference, rather than fixed generation in advance. As a result, conducting experiments on GPUs may lead to slower overall inference speeds. Extending this paradigm to brain-inspired chips with asynchronous event input is an interesting direction worth exploring in the future.
@@ I.3 Details of Dynamic Hyperparameter Tuning
Appendix J Visualization of the SNN Training Process
To validate whether the proposed feedback-update strategy can serve a guiding role during the initial stages of training, we have visualized the training process of the SNN and presented it in Figure 1 of the supplementary PDF rebuttal file. As anticipated, the training of the SNN exhibited fluctuations during the initial training stage, which might be attributed to the instability of the event quality obtained from dynamic slicing at this early phase. However, as training progressed, the loss of the SNN gradually stabilized and decreased, converging towards a desired outcome. Correspondingly, the slicing times progressively converged towards the desired spiking index. Therefore, although initial exploration may require several steps, our proposed training method is capable of offering effective guidance.
Figure 10: Visualization of the SNN Training Process. As shown in the gray box in (a), the SNN loss fluctuates considerably during the initial training stage. This is due to the instability in the quality of dynamic sliced events during the early phase (warm-up stage). However, as training progresses, the SNN loss gradually stabilizes and converges (guidance stage), with the corresponding slicing time (b) also converging towards the desired spike index.
Appendix K Implementation Details in Toy Experiments
@@
In detail, we set the time step within the range $[1,30]$ and the max number of iterations as 800. We utilize a lightweight convolutional SNN with random initialization. We compare our proposed SPA-Loss function with common mean square error (MSE-Loss) and cross-entropy loss (CE-Loss). The results in Fig. 4 demonstrate that our proposed SPA-Loss can guide the SNN to spike at the desired timestep accurately with fewer convergence iterations than standard MSE-Loss and CE-Loss, paving the way for complex event-based vision tasks in Sec. 4.2. More experiments on the beginner’s arena are shown in Appendix N.
Appendix L Implementation Details in Event-based Task
@@
L.1 Event-based Object Tracking
@@
L.2 Event-based Recognition
@@
Appendix M Theoretical Energy Consumption Calculation
@@
Appendix N More Experiments on Beginner’s Arena
@@
Appendix O Statistics of Dynamic Slicing (SpikeSlicer) vs. Fixed Slicing
@@
Appendix P Statistics of Resulting Slicing Duration
@@
Appendix Q More Experiments with Latest Models
@@
Appendix R More Experiments with Complex Neuromorphic Dataset
@@
Appendix S Impact Statement